# via Dodecaneso 35, 16146, Genova, Italy

### v1

@@ Abstract
Multivariate Cryptography is one of the main candidates for Post-quantum Cryptography. Multivariate schemes are usually constructed by applying two secret affine invertible transformations $\mathcal{S},\mathcal{T}$ to a set of multivariate polynomials $\mathcal{F}$ (often quadratic). The secret polynomials $\mathcal{F}$ posses a trapdoor that allows the legitimate user to find a solution of the corresponding system, while the public polynomials $\mathcal{G}=\mathcal{S}\circ\mathcal{F}\circ\mathcal{T}$ look like random polynomials. The polynomials $\mathcal{G}$ and $\mathcal{F}$ are said to be affine equivalent. In this article, we present a more general way of constructing a multivariate scheme by considering the CCZ equivalence, which has been introduced and studied in the context of vectorial Boolean functions.
@@ 1 Introduction
A multivariate public-key cryptosystem involves a public key comprising multivariate polynomials $f^{(1)},\ldots,f^{(m)}$ in ${\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$, where ${\mathbb{F}}_{q}$ is a finite field with $q$ elements. In order to keep the public key size not too large, usually, quadratic polynomials are considered. The secret key is some information (about the construction of the polynomials $f^{(i)}$) which allows to solve the system $f^{(1)}=y_{1},\ldots,f^{(m)}=y_{m}$ efficiently for some $y_{1},\ldots,y_{m}\in{\mathbb{F}}_{q}$. To encrypt a message $(x_{1}^{\prime},\ldots,x_{n}^{\prime})\in{\mathbb{F}}_{q}^{n}$, the sender computes $y_{i}=f^{(i)}(x_{1}^{\prime},\ldots,x_{n}^{\prime})$, for $i=1,\ldots,m$, and sends $(y_{1},\ldots,y_{m})$ to the receiver. With the secret key, the receiver can solve the system and recover the original message.
One of the main methods to achieve the previous scheme is the Bipolar Construction. From a system $\mathcal{F}$ of $m$ equations in $n$ variables (possibly quadratic) relatively easy to invert, the secret key is composed of $\mathcal{F}$ and two randomly chosen affine bijections $\mathcal{S}$ and $\mathcal{T}$; the public key is the system $\mathcal{G}=\mathcal{S}\circ\mathcal{F}\circ\mathcal{T}$. The obtained system $\mathcal{G}$ (still quadratic) is now assumed to be not easy to invert, since it should be hardly distinguishable from a random system. Notice that, in this case, finding a preimage of $y=(y_{1},\ldots,y_{m})$ reduces to finding a preimage for $\mathcal{F}$ of $\mathcal{S}^{-1}(y)$ and then apply $\mathcal{T}^{-1}$. Hence, the core idea of the Bipolar Construction is to hide the structure of the central map $\mathcal{F}$ by applying two random affine bijections to the input and to the output of $\mathcal{F}$. This corresponds to randomly taking a system in the affine-equivalence class of $\mathcal{F}$.
@@
In order to better explain this phenomenon, we briefly recall the example of the Matsumoto–Imai (MI) Cryptosystem [MI88]. In this scheme, the authors consider $q$ a power of two and $m=n$. The central system $\mathcal{F}$ is seen as a quadratic function $F:\mathbb{E}\rightarrow\mathbb{E}$, $F(x)=x^{q^{j}+1}$ where $\mathbb{E}=\mathbb{F}_{q^{n}}$ and $j$ is chosen such that $\gcd(q^{n}-1,q^{j}+1)=1$. Then, by applying the standard isomorphism $\phi:{\mathbb{F}}_{q}^{n}\rightarrow\mathbb{E}$, the map is expanded as a system of $n$ equations over ${\mathbb{F}}_{q}$, $\mathcal{F}=\phi\circ F\circ\phi^{-1}:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{n}$. The quadratic function $F(x)=x^{q^{j}+1}$ is a bijection and it is easy to invert. Clearly, after applying the affine bijections $\mathcal{S}$ and $\mathcal{T}$ to the input and to the output of $\mathcal{F}$, the system will not have a so-simple structure easy to invert. However, the function $F$ presents a linear relation between its input and its output. Indeed, setting $y=F(x)$, we have the relation $y^{q^{j}}x=x^{q^{2j}}y$, in which both variables $x$ and $y$ appear only with “linear exponents”. The affine transformations $\mathcal{S}$ and $\mathcal{T}$ preserve the presence of such linear relations between input and output, making systems susceptible to linearization attacks as first demonstrated in [Pat95]. Modifications to the MI cryptosystem have been attempted to mitigate this vulnerability, as seen in [CGP92, Din04]. However, even with these alterations, attacks have been developed that can reduce the modified MI system back to its original form, thus rendering it vulnerable to linearization attacks, as detailed in [DFSS07, FGS05]. This exemplifies just one instance of a property that persists through an affine equivalence relation. Cryptographers have identified many such properties, which have been exploited in successful attacks on multivariate schemes, including the Min Rank attack documented in [BFP13, CG21a, LC00, KS98, VST17].
@@
In this paper, we propose to use a more general notion of equivalence relation between polynomial systems to obtain a public-key function $\mathcal{G}$ which does not inherit the “simple” structures of the secret function $\mathcal{F}$. Specifically, our investigation focuses on the CCZ equivalence transformation which has been introduced by Carlet, Charpin, and Zinoviev [CCZ98]. Given two functions $F,G:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$ we say that they are CCZ equivalent if there exists an affine bijection $\mathcal{A}$ of ${\mathbb{F}}_{q}^{n+m}$ such that $\mathcal{G}_{G}=\mathcal{A}(\mathcal{G}_{F})$, where $\mathcal{G}_{F}$ and $\mathcal{G}_{G}$ are the graphs of $F$ and $G$ respectively. This equivalence relation has been mainly studied in the context of cryptographic Boolean functions, since it keeps unchanged the value of the differential uniformity and the nonlinearity, two important properties to study when a function is used as a component of a block cipher. Budaghyan, Carlet, and Pott proved that CCZ equivalence is strictly more general than the affine equivalence by exhibiting functions which are CCZ equivalent to $F(x)=x^{3}$ over $\mathbb{F}_{2^{n}}$ , but not affine equivalent, and also not extended affine (EA) equivalent [BCP06]. Other works focused on further studying the relation between CCZ and EA equivalences, see for example [CP19, BCV20].
@@
We propose constructing multivariate schemes where the secret map and the public map are CCZ equivalent, without necessarily being affine or extended affine equivalent. However, selecting a random element $G$ in the CCZ class of a given polynomial map $F$ presents a challenge. Not every affine bijection of ${\mathbb{F}}_{q}^{n+m}$ maps the graph of one function into the graph of another, and the admissible affine bijections depend on the chosen function $F$. To address this challenge, we leverage a result by Canteaut and Perrin [CP19], demonstrating that any two CCZ-equivalent functions can be connected by applying two extended affine transformations and another map called a $t$-twist (see Definition 3.1). In Section 4, we provide a high-level explanation of how this strategy can be utilized to construct an encryption or a signature scheme. This construction is quite general and applicable to any secret map $F$ that admits a $t$-twist. To provide a concrete example, we propose selecting a quadratic function $F$ derived from Oil and Vinegar (OV) polynomials [Pat95]. OV polynomials divide variables into oil and vinegar sets, with no quadratic terms in the oil variables, allowing for linear decryption/signing by assigning random values to the vinegar variables. The name comes from the fact that the variables do not truly mix, like oil and vinegar in the salad dressing. We name our scheme Pesto, as the CCZ transformation ensures that the variables fully mix, resembling the mixing of ingredients in Pesto Sauce using mortar and pestle. Notably, while the secret polynomials are quadratic, the public polynomials have degree four.
@@
Since our primary objective is to establish a connection between the research areas of Boolean functions and Multivariate Cryptography rather than presenting a fully-fledged new proposal, we have opted not to propose specific parameters for the system. We believe that there is considerable scope for further research and development in this area.
@@
This work is organized as follows. In Section 2 we recall the definition and some basic facts about the CCZ equivalence. In Section 3, we present the twisting and explain how it can be used to produce a random CCZ transformation. In Section 4 we present a proposal for a multivariate cryptographic scheme obtained by hiding the central map with a CCZ transformation. We present it at a high level of generality in Subsection 4.1, and then with more specifics on the functions in Subsection 4.2, i.e., the Pesto scheme. Section 5 presents some comments on the form of the constructed maps, an analysis on the sizes of the keys and on the computational cost of applying the proposed procedure. In the last section, we present a preliminary analysis on the security of the scheme.
@@ Acknowledgments
A. Caminata and I. Villa are supported by the Italian PRIN2022 grant P2022J4HRR “Mathematical Primitives for Post Quantum Digital Signatures”, by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001, and by the European Union within the program NextGenerationEU. Additionally, part of the work was done while Caminata was visiting the Institute of Mathematics of the University of Barcelona (IMUB). He gratefully appreciates their hospitality during his visit.
@@ 2 Preliminaries
In this section we recall the definition of CCZ equivalence, introduced in [CCZ98], together with some preliminary results and notations.
@@ 2.1 CCZ equivalence
Let $n,m$ be positive integers, $q$ a prime power and ${\mathbb{F}}_{q}$ a finite field with $q$ elements. We consider a function $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$. Notice that $F$ can be seen as $F=(f_{1},\ldots,f_{m})$ with $f_{i}:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}$ for $1\leq i\leq m$. We call coordinate, or $i$-th coordinate, of $F$ the function $f_{i}$. For $\lambda=(\lambda_{1},\ldots,\lambda_{m})\in{\mathbb{F}}_{q}^{m}$ we call the $\lambda$-component of $F$ the function $F_{\lambda}=\lambda\cdot F=\lambda_{1}f_{1}+\cdots+\lambda_{m}f_{m}$. To represent $F$ we can use the algebraic normal form (ANF), that is, we can represent the function as a multivariate polynomial over ${\mathbb{F}}_{q}^{m}$:
$F(x)=F(x_{1},\ldots,x_{n})=\sum_{u\in\mathbb{N}^{n}}a_{u}x_{1}^{u_{1}}\cdots x_{n}^{u_{n}},\qquad\mbox{ with }a_{u}\in{\mathbb{F}}_{q}^{m}.$
Moreover, in order to have a unique representative for $F$ we adopt the standard convention that $u_{1},\dots,u_{n}<q$. We say that $x_{1}^{u_{1}}\cdots x_{n}^{u_{n}}$ is a term of $F$ if $a_{u}\neq 0$. The algebraic degree of $F$ is $\deg(F)=\max\{\sum_{i=1}^{n}u_{i}\,:\,u\in\mathbb{N}^{n};\,a_{u}\neq 0\}$. We call $F$ linear if $\deg(F)=1$ and $F(0)=0$, affine if $\deg(F)\leq 1$, quadratic if $\deg(F)\leq 2$. When $m=n$, we say that $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{n}$ is a bijection or that it is invertible if $F$ induces a permutation over ${\mathbb{F}}_{q}^{n}$ , i.e.,  if $\{F(v):v\in{\mathbb{F}}_{q}^{n}\}={\mathbb{F}}_{q}^{n}$.
@@
  • $F$ is CCZ equivalent to $G$ if there exists an affine bijection $\mathcal{A}$ of ${\mathbb{F}}_{q}^{n+m}$ such that $\mathcal{G}_{G}=\mathcal{A}(\mathcal{G}_{F})$, where $\mathcal{G}_{F}=\{(x,F(x)):x\in{\mathbb{F}}_{q}^{n}\}\subseteq{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ is the graph of $F$, and $\mathcal{G}_{G}$ is the graph of $G$.
@@
In the usual set up of multivariate schemes, the central (secret) map $F$ is composed with two randomly chosen affine bijections $A_{1},A_{2}$ of ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively to obtain the public map $G=A_{1}\circ F\circ A_{2}$. Thus, the secret and public key $F$ and $G$ are affine equivalent. Modifying this construction by using EA equivalence does not provide any improvement. Indeed, the difference between EA and affine equivalence is just the addition of an affine transformation. So, most of the attacks that can be performed over schemes using an affine transformation can be easily extended to the case of EA transformation. Therefore, our investigation will focus on the CCZ transformation. Unfortunately, given a function $F$, it seems not so easy to obtain a random CZZ-equivalent function $G$ as we are going to explain next.

### v2

@@ Abstract
Multivariate Cryptography is one of the candidates for Post-quantum Cryptography. Multivariate schemes are usually constructed by applying two secret affine invertible transformations $\mathcal{S},\mathcal{T}$ to a set of multivariate polynomials $\mathcal{F}$ (often quadratic). The polynomials $\mathcal{F}$ possess a trapdoor that allows the legitimate user to find a solution of the corresponding system, while the public polynomials $\mathcal{G}=\mathcal{S}\circ\mathcal{F}\circ\mathcal{T}$ look like random polynomials. The polynomials $\mathcal{G}$ and $\mathcal{F}$ are said to be affine equivalent. In this article, we present a more general way of constructing a multivariate scheme by considering the CCZ equivalence, which has been introduced and studied in the context of vectorial Boolean functions.
@@ 1 Introduction
A multivariate public-key cryptosystem involves a public key comprising multivariate polynomials $f^{(1)},\ldots,f^{(m)}$ in ${\mathbb{F}}_{q}[z_{1},\ldots,z_{n}]$, where ${\mathbb{F}}_{q}$ is a finite field with $q$ elements. In order to keep the public key size not too large, usually, quadratic polynomials are considered. The secret key is some information (about the construction of the polynomials $f^{(i)}$) which allows to solve the system $f^{(1)}=w_{1},\ldots,f^{(m)}=w_{m}$ efficiently for some $w_{1},\ldots,w_{m}\in{\mathbb{F}}_{q}$. To encrypt a message $(z_{1}^{\prime},\ldots,z_{n}^{\prime})\in{\mathbb{F}}_{q}^{n}$, the sender computes $w_{i}=f^{(i)}(z_{1}^{\prime},\ldots,z_{n}^{\prime})$, for $i=1,\ldots,m$, and sends $(w_{1},\ldots,w_{m})$ to the receiver. With the secret key, the receiver can solve the system and recover the original message.
One of the main methods to achieve the previous scheme is the Bipolar Construction. From a system $\mathcal{F}$ of $m$ equations in $n$ variables (possibly quadratic) relatively easy to invert, the secret key is composed of $\mathcal{F}$ and two randomly chosen affine bijections $\mathcal{S}$ and $\mathcal{T}$; the public key is the system $\mathcal{G}=\mathcal{S}\circ\mathcal{F}\circ\mathcal{T}$. The obtained system $\mathcal{G}$ (still quadratic) is now assumed to be not easy to invert, since it should be hard to distinguish from a random system. Notice that, in this case, finding a preimage of $(w_{1},\ldots,w_{m})$ reduces to finding a preimage for $\mathcal{F}$ of $\mathcal{S}^{-1}(w_{1},\ldots,w_{m})$ and then apply $\mathcal{T}^{-1}$. Hence, the core idea of the Bipolar Construction is to hide the structure of the central map $\mathcal{F}$ by applying two random affine bijections to the input and to the output of $\mathcal{F}$. This corresponds to randomly taking a system in the affine-equivalence class of $\mathcal{F}$.
@@
In order to better explain this phenomenon, we briefly recall the example of the Matsumoto–Imai (MI) Cryptosystem [MI88]. In this scheme, the authors consider $q$ a power of two and $m=n$. The central system $\mathcal{F}$ is seen as a quadratic function $F:\mathbb{E}\rightarrow\mathbb{E}$, $F(z)=z^{q^{j}+1}$ where $\mathbb{E}=\mathbb{F}_{q^{n}}$ and $j$ is chosen such that $\gcd(q^{n}-1,q^{j}+1)=1$. Recall that the degree of a monomial $z^{k}\in\mathbb{E}[z]$, $10\leq k\leq q^{n}-$ 1, is the 2-weight of $k$. Then, by applying the standard isomorphism $\phi:{\mathbb{F}}_{q}^{n}\rightarrow\mathbb{E}$, the map is expanded as a system of $n$ equations over ${\mathbb{F}}_{q}$, $\mathcal{F}=\phi\circ F\circ\phi^{-1}:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{n}$. The quadratic function $F(z)=z^{q^{j}+1}$ is a bijection and it is easy to invert. Clearly, after applying the affine bijections $\mathcal{S}$ and $\mathcal{T}$ to the input and to the output of $\mathcal{F}$, the system will not have a so-simple structure easy to invert. However, the function $F$ presents a bilinear relation between its input and its output. Indeed, setting $w=F(z)$, we have the relation $w^{q^{j}}z=z^{q^{2j}}w$, in which both variables $z$ and $w$ appear only with “linear exponents”. The affine transformations $\mathcal{S}$ and $\mathcal{T}$ preserve the presence of such linear relations between input and output, making systems susceptible to linearization attacks as first demonstrated in [Pat95]. Modifications to the MI cryptosystem have been attempted to mitigate this vulnerability, as seen in [CGP92, Din04]. However, even with these alterations, attacks have been developed that can reduce the modified MI system back to its original form, thus rendering it vulnerable to linearization attacks, as detailed in [DFSS07, FGS05, ØFR24]. This exemplifies just one instance of a property that persists through an affine equivalence relation. Cryptographers have identified many such properties, which have been exploited in successful attacks on multivariate schemes, including the Min Rank attack documented in [BFP13, CG21a, LC00, KS98, VST17].
@@
In this paper, we propose to use a more general notion of equivalence relation between polynomial systems to obtain a public-key function $\mathcal{G}$ which does not inherit the “simple” structures of the secret function $\mathcal{F}$. Specifically, our investigation focuses on the CCZ equivalence transformation which has been introduced by Carlet, Charpin, and Zinoviev [CCZ98]. Given two functions $F,G:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$ we say that they are CCZ equivalent if there exists an affine bijection $\mathcal{A}$ of ${\mathbb{F}}_{q}^{n+m}$ such that $\mathcal{G}_{G}=\mathcal{A}(\mathcal{G}_{F})$, where $\mathcal{G}_{F}$ and $\mathcal{G}_{G}$ are the graphs of $F$ and $G$ respectively. This equivalence relation has been mainly studied in the context of cryptographic Boolean functions, since it keeps unchanged the value of the differential uniformity and the nonlinearity, two important properties to study when a function is used as a component of a block cipher. Budaghyan, Carlet, and Pott proved that CCZ equivalence is strictly more general than the affine equivalence by exhibiting functions which are CCZ equivalent to $F(z)=z^{3}$ over $\mathbb{F}_{2^{n}}$, but not affine equivalent, and also not extended affine (EA) equivalent [BCP06], where EA equivalence extends affine equivalence by adding any affine transformation. Other works focused on further studying the relation between CCZ and EA equivalences, see for example [CP19, BCV20]. In this work, we propose constructing multivariate schemes in which the secret and public maps are CCZ equivalent but not necessarily affine or extended affine equivalent. While CCZ equivalence preserves the cryptographic properties mentioned above, these may not be relevant in this context. Instead, we focus on the fact that CCZ equivalence is more general than affine equivalence and does not preserve other key properties such as algebraic degree and bijectivity. The central idea of this paper is to investigate whether the relations usually used to break multivariate systems—typically preserved under affine transformations—are instead disrupted by CCZ transformations.
@@
One of the main challenges in transitioning from affine equivalence to CCZ equivalence is that selecting a random element $G$ from the CCZ class of a given polynomial map $F$ is not straightforward. Not every affine bijection of ${\mathbb{F}}_{q}^{n+m}$ maps the graph of one function into the graph of another, and the admissible affine bijections depend on the chosen function $F$. To address this challenge, we leverage a result by Canteaut and Perrin [CP19], demonstrating that any two CCZ equivalent functions can be connected by applying two extended affine transformations and another map called a $t$-twist (see Definition 3.1). In Section 4, we provide a high-level explanation of how this strategy can be utilized to construct an encryption or a signature scheme. This construction is quite general and applicable to any secret map $F$ that admits a $t$-twist. To provide a concrete example, we propose selecting a quadratic function $F$ derived from Oil and Vinegar (OV) polynomials [Pat95]. OV polynomials divide variables into oil and vinegar sets, with no quadratic terms in the oil variables, allowing for linear decryption/signing by assigning random values to the vinegar variables. The name comes from the fact that the variables do not truly mix, like oil and vinegar in the salad dressing. We name our scheme Pesto, as the CCZ transformation ensures that the variables fully mix, resembling the mixing of ingredients in Pesto Sauce using mortar and pestle. Notably, while the secret polynomials are quadratic, the public polynomials have degree four.
@@
Although our primary objective is to connect the research areas of Boolean functions and Multivariate Cryptography rather than presenting a fully developed proposal, we also explore specific parameter choices for the system. These parameters are not intended for immediate implementation but rather serve as an invitation for further cryptanalysis of the scheme. We believe this area offers significant opportunities for further research and development, and we warmly encourage additional contributions.
@@
This work is organized as follows. In Section 2 we recall the definition and some basic facts about the CCZ equivalence. In Section 3, we present the twisting and explain how it can be used to produce a random CCZ transformation. In Section 4 we present a proposal for a multivariate cryptographic scheme obtained by hiding the central map with a CCZ transformation. We present it at a high level of generality in Subsection 4.1, and then with more specifics on the functions in Subsection 4.2, i.e., the Pesto scheme. Section 5 presents some comments on the form of the constructed maps, an analysis on the sizes of the keys and on the computational cost of applying the proposed procedure. In the last section, we present a preliminary analysis on the security of the scheme and propose a possible choice of secure parameters.
@@ Acknowledgments
A. Caminata is supported by the Italian PRIN2022 grant P2022J4HRR “Mathematical Primitives for Post Quantum Digital Signatures”, by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001, and by the European Union within the program NextGenerationEU. Additionally, part of the work was done while Caminata was visiting the Institute of Mathematics of the University of Barcelona (IMUB). He gratefully appreciates their hospitality during his visit.
The research of M. Calderini and I. Villa was partially supported by the Italian Ministry of University and Research with the project PRIN 2022RFAZCJ.
The three authors are members of the INdAM Research Group GNSAGA.
We sincerely thank the anonymous referees for their insightful comments and for encouraging us to propose specific parameters for the scheme.
@@ 2 Preliminaries
In this section, we recall the definition of CCZ equivalence, introduced in [CCZ98], together with some preliminary results and notations.
@@ 2.1 CCZ equivalence
Let $n,m$ be positive integers, $q$ a prime power and ${\mathbb{F}}_{q}$ a finite field with $q$ elements. We consider a function $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$. Notice that $F$ can be seen as $F=(f_{1},\ldots,f_{m})$ with $f_{i}:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}$ for $1\leq i\leq m$. We call the coordinate, or the the $i$-th coordinate, of $F$ the function $f_{i}$. For $\lambda=(\lambda_{1},\ldots,\lambda_{m})\in{\mathbb{F}}_{q}^{m}$ we call the $\lambda$-component of $F$ the function $F_{\lambda}=\lambda\cdot F=\lambda_{1}f_{1}+\cdots+\lambda_{m}f_{m}$. To represent $F$ we can use the algebraic normal form (ANF), that is, we can represent the function as a multivariate polynomial over ${\mathbb{F}}_{q}^{m}$:
$F(z)=F(z_{1},\ldots,z_{n})=\sum_{u\in\mathbb{N}^{n}}a_{u}z_{1}^{u_{1}}\cdots z_{n}^{u_{n}},\qquad\mbox{ with }a_{u}\in{\mathbb{F}}_{q}^{m}.$
Moreover, in order to have a unique representative for $F$ we adopt the standard convention that $u_{1},\dots,u_{n}<q$. We say that $z_{1}^{u_{1}}\cdots z_{n}^{u_{n}}$ is a term of $F$ if $a_{u}\neq 0$. The algebraic degree of $F$ is $\deg(F)=\max\{\sum_{i=1}^{n}u_{i}\,:\,u\in\mathbb{N}^{n};\,a_{u}\neq 0\}$. We call $F$ linear if $\deg(F)=1$ and $F(0)=0$, affine if $\deg(F)=1$, quadratic if $\deg(F)=2$. When $m=n$, we say that $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{n}$ is a bijection or that it is invertible if $F$ induces a permutation over ${\mathbb{F}}_{q}^{n}$ , i.e.,  if $\{F(v):v\in{\mathbb{F}}_{q}^{n}\}={\mathbb{F}}_{q}^{n}$.
@@
  • $F$ is CCZ equivalent to $G$ if there exists an affine bijection $\mathcal{A}$ of ${\mathbb{F}}_{q}^{n+m}$ such that $\mathcal{G}_{G}=\mathcal{A}(\mathcal{G}_{F})$, where $\mathcal{G}_{F}=\{(v,F(v)):v\in{\mathbb{F}}_{q}^{n}\}\subseteq{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ is the graph of $F$, and $\mathcal{G}_{G}$ is the graph of $G$.
@@
In the usual setup of multivariate schemes, the central (secret) map $F$ is composed with two randomly chosen affine bijections $A_{1},A_{2}$ of ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively to obtain the public map $G=A_{1}\circ F\circ A_{2}$. Thus, the secret and public key $F$ and $G$ are affine equivalent. Modifying this construction by using EA equivalence does not provide any improvement. Indeed, the difference between EA and affine equivalence is just the addition of an affine transformation. So, most of the attacks that can be performed over schemes using an affine transformation can be easily extended to the case of EA transformation. Therefore, our investigation will focus on the CCZ transformation. Unfortunately, given a function $F$, it seems not so easy to obtain a random CCZ equivalent function $G$ as we are going to explain next.